ABCDEFGHIJKLMNOPQRSTUVWXYZAAABAC
1
run/timeRun originatingSuperceded?typeValiditycommentslogbcm_an_us ped
bcm_an_us scale
150 signalped percentage-2.047925911
2
PREX Version: https://docs.google.com/spreadsheets/d/1n-zEq-t44vxhzQOG_hioxDPO0-x1MJo53fyWXbst7qo/edit#gid=0
All until proven otherwise-288.540.009164216368.04085-1.7628255120.2851003992
3
5396Good for 50-100unser50-100 uA: runs 5396->54194 points, beam trips, 150uA
https://logbooks.jlab.org/entry/3753060
-606.310.00891116833.1276-3.601885607-1.553959695
4
5410PQB health check injtest, not a calibration run: PQB health check - injector
https://logbooks.jlab.org/entry/3753030
PQB health check - injector
5
5416Aq issuesBad run
large bcm asym means => aq feedback might have been off. Deemed bad
6
54175417yesdetector/sam5410-542030-130uA, 4 good points, sams with different preamps than later
https://logbooks.jlab.org/entry/3753970
7
5420 or so, 12/13/19 - morning5421event:SAM preamps changed during accessallSAM4, SAM3 and SAM7 preAmps were changed
https://logbooks.jlab.org/entry/3752894
8
12/13/2019, ~10am, run54225421 (junk)event: DN quartz switched to counting
https://logbooks.jlab.org/entry/3752894
9
12/13/19 DAY 5424 (junk)
event:FFB commissioning and compton rates tuning straight through
all
https://logbooks.jlab.org/entry/3753087
10
5433BPMAllno FFB, iocse10 and iocse9 gains off at 143uA, 110-145uA 5points
https://logbooks.jlab.org/entry/3753602
11
54345434event:SAMs 3,7 HV changedall
12
5434unserAll, except cav4cQ, 5420->54748 points, FFB off, 150 to 30
https://logbooks.jlab.org/entry/3753564
-5970.00897516713.09192-3.57205-1.524124089
13
5435event: SAM HV changedHigh gain SAM HV was changed during this run
https://logbooks.jlab.org/entry/3753239
14
5436detector/sam
All, except SAM 7 later, use 100-150 fit range
2x2 raster (1.8x1.4MCC), FFB on, decreasing currents, gain searches done at 70,50,30uA
https://logbooks.jlab.org/entry/3753968
15
5438PITA Scan, RHWP 1067??
https://prex.jlab.org/wiki/index.php/WAC_Notes_December_14_2019
16
12/14/2019 ~15:15 DetectorMAIN detectors moved just before 5454
https://logbooks.jlab.org/entry/3753596
17
12/14/2019 18:37:00 run5456event: main detector moveall
https://logbooks.jlab.org/entry/3753596
18
12/14 DAY
event:FFB ioc lost heartbeat (afternoon?), SL E corrected manually 12X pushed from ~0.5mm toward 0mm, HALLA:dpp changed by -0.0002 towards 0, apparently the FFB choice of bpm selection was changed according to Yves
all
https://logbooks.jlab.org/entry/3753542
19
12/15/2019 during run5456, run5457
AT moveATs were moved
https://logbooks.jlab.org/entry/3753667
20
12/14/2019 during run5460AT move and HV resetATs again moved and reset HV
https://logbooks.jlab.org/entry/3753667
21
5460event:SAM7 HV adjustment, AT HV adjustmentall
22
5473
event:6dB cav4c attenuation added on inpputs to dig reciever
all
https://logbooks.jlab.org/entry/3753874
23
5473
event: starting with this run dithering ampltudes increased by 2x
https://logbooks.jlab.org/entry/3753871
24
5474event: start of run SAM HV adjustment
From run 5474 -> for cav4cQ and ATs and SAM7
https://logbooks.jlab.org/entry/3753866
25
5474detector/samFrom 5474 -> for ATs and SAMs
FFB off, 20-150uA, raster 2x2 (1.8x1.4MCC), increasing currents, X=-0.3,Y=-0.5mm
https://logbooks.jlab.org/entry/3753933
26
12/15/2019 just before run5474SAM7SAM7 HV lowered
https://logbooks.jlab.org/entry/3753866
27
5475ProductionFFB Finally works
https://prex.jlab.org/wiki/index.php/WAC_Notes_December_15_2019
28
12/15/2019 22:50 run5478AT move
https://logbooks.jlab.org/entry/3754116
29
5479
event: starting with this run dithering ampltudes increased
https://logbooks.jlab.org/entry/3753955
30
31
32
12/16/2019 before run5483AT moveATin moved in LHRS
https://logbooks.jlab.org/entry/3754929
33
1/1-1/3
cathode reheated, rotated 180degrees, fine tune rotation, new rhwp setting, is settings, pita
34
1/4/2020-1/5/2020
attenuators on 4B and 4D,cav4b,4d split signal to Yury
4B and 4D I 20dB atten removed. splitter inserted (3dB) for I. 6dB atten added on dig recievers for 4B and 4D XI and YI muxed.
https://logbooks.jlab.org/entry/3758879
35
1/6/2020 run5554
cav4B,4D digital internal gain set from VGA1=6, VGA2=31 to VGA1=0,VGA2=0. to match cav4c settings which work at high current
log 3759230
36
1/7/2020 6:30 AM
usl base pmt replaced after counting mode measurements done with 2000V
37
1/7/2020Vwien, solenoid fine tune
38
1/7/2020 7:00 PMspin dance, HWien new =-29.64
39
run 5566unser like calibration (50-80 uA)
40
5572
hel mag Horiz 0l03H 6000,-6000 0l02H -3000,3000
https://logbooks.jlab.org/entry/3759880
41
5574
hel mag very 0l01V 6000,-6000 0l03V -3000,3000
https://logbooks.jlab.org/entry/3759883
42
1/8 early morning run5576unser like current ramp with beam trips
https://logbooks.jlab.org/entry/3760623
-645.630.00892416808.60601-3.84106808-1.793142169
43
run 5576 1/8
slow ramp up on Ca48 2x2 raster, like set calibration
same run doubles as both calibrations
44
1/8 afternoon straight through Compton rate tuning Yves
45
1/8 swing 5584
cav4B,d DACSL1,2 changed from 0x0 0x0 to 0x1 0x1=0x101.
4b and 4d have no signal in parity daq up til this point
46
1/9 owl, run5590-5595
quad 1H01 adjust, affects BMW slopes, reg slopes, spot size, bpm widths
47
1/9 swing run 5599Yes
cav 4B,4D add 6dB atten so total atten 18dB on rec. input
48
1/10 day run56185618pita scan 5618
49
run5626, before, swing 1/10
left dipole and quad slight corrections, likely negligable
https://logbooks.jlab.org/entry/3761369
50
1/10 16:38Yescav4B,cav4D 6dB atten added. 24dB atten total
https://logbooks.jlab.org/entry/3761328
51
1/11
FFB turning on now causes trips/jumps in beam positions all around the accelerator -> OPS has an issue with re-establishing injector/orbit locks. Fixes itself much later in February
https://logbooks.jlab.org/entry/3761789
52
1/11/2020 owl, slug 106
detector widths increased, yield lower, so rates lower, discovered left arm alignment changed (gausmeter not reliable measure, perhaps related to ioc communication error Wed/Thursday
https://logbooks.jlab.org/entry/3761778
53
1/11 10:005636Yes, at run 5645
detector alignment check and adjustment on left arm with dipole current
https://logbooks.jlab.org/entry/3761792
54
1/11 day run5646LQ1 tripped, rarm running only
55
1/11 day run5647cycled LQ1 2arm running again
56
1/11 swing start approx before run 5655
FFB Fixed
57
58
1/11 swing run56535653
calibration: bpm with manual gains on ffb , gains off for 150-120uA, then autogain for 100-20uA, detector, bcm, cavity, bpm calibration all in one
59
1/13 dayRun 5691
Changing attenuators in the hall and in the CH to get Cavities B and D to work
28 dD added to B/D I channels in Hall, 24->down to 6 dB in CH
https://logbooks.jlab.org/entry/3763028
60
1/15 swing
tune through Compton chicane, optics change of dispersion in ARC
https://logbooks.jlab.org/entry/3784872
61
1/16 owl
1/16/2020 owl run5744-5746
BMW tune 1H01, raster change to 1.5x2 MCC
62
5750Compton tuning
Changes BPM1x down by 3mm, RHRS asym grows, slopes of 4eX and 12X change a lot
https://prex.jlab.org/wiki/index.php/WAC_Notes_January_16_2020
63
1/16 Swing5761ATout moved in RHRS
https://logbooks.jlab.org/entry/3765195
64
1/17 day 8:30-9amrun 5778
calibration, bpms from 120-150uA (scandata2=1,2 bad), everything else AT's cav4B,D, sams, sets, bcms, can have data points 20-150uA
65
1/18 23:41run5819
target death, ops changes MVS1P04V with lock and manually fatally moving beam position by 2mm without tgt orbit lock enabled
http://ace.phys.virginia.edu/HAPPEX/4010
66
1/20 day5849at IN both arms moved, physically
https://logbooks.jlab.org/entry/3767677
67
1/21targets, Ca40 in mineral oil bath
68
1/22 day 4:20run5852
1p02b XM cable unplugged, plugged in readout 10X higher, needs new pedestal. runs before this need a scale factor applied to the XM channel based on calibration above or disregard bpm 1p02b alroghter.l
69
1/23
cav4B,d examined downstairs, it was decided to change VGA gains later when beam is back....
https://logbooks.jlab.org/entry/3768582
70
1/24 owlrun5865detectors moved before this run (MAIN plus AT)
https://logbooks.jlab.org/entry/3769574
https://logbooks.jlab.org/entry/3769736
71
1/25 owlrun5897detectors moved
https://logbooks.jlab.org/entry/3770283
72
1/25 owl-dayslug115
beam position lowered -0.3,-1.5 and new Ca48 target, not through compton, only 120uA
73
1/25 swingrun5924
detectors moved, just before this run, to minimize inelastic state contamination
https://logbooks.jlab.org/entry/3770683
https://logbooks.jlab.org/entry/3770808
74
1/25 swing -1/27 dayslug117-slug120AposU increased by 400 starting next slug117
https://logbooks.jlab.org/entry/3770772
75
1/27 day -1/27 dayMoller measurement, AposU +400 from nominal
76
1/27 swing, slug121run5991??
AposU decreased by 400 from nominal, Moller left Ihwp out, hopefully slug 120 got incremented to slug121 at this time
77
1/29 day6035Major DAQ synch loss. Experts come in and fix
78
1/31 before 9:30 am, slug122?,123,124
slug123 , Ihwpout,will have AposU+400 from nominal, slug124 ihwpin will have AposV+400 from nominal. flip values macro edited in morning...not used until next slug.
79
2/1 10:30pm slugs125,126
flipper edited, slugs 25-out will have AposV+400 from nominal,and 26-in to have AposV-400 from nominally
probably wrong
80
2/2 8:20 am slugs126,127
flipper edited, slugs 27-out will have AposV-400 from nominal,and 26-in to have AposV-400 from nominal
probably wrong
81
2/2 9:25pm slugs128,129
flipper not edited, slugs 29-out will have Aposv-400 from nominal,and 28-in to have Aposv-400 from nominal
probably wrong
82
2/2 11pm ish
1H01 changed for beam mod,. not going through chicane due to rates, 1c16 quad changed by ops, have been using Compton lock with one bpm instead of 2.
83
2/3 6:30-8pm6176
pseudo-unser like calibration (beam trippy), also current ramp-like since ffb on for all currents and slow ramp up waiting for beam positions to stabilize. x,y=-0.3,-1.5. raster 1.7x2.6 MCC .
applies to run5897 onwards, could be used on runs after new target installed 5853ish
Caryn, Amali analyzed
https://logbooks.jlab.org/entry/3777612
235.090.00955815693.659761.497993483.545919391
84
2/4production running at 140uAsome 150uA production runs during day, then to 140uA
https://logbooks.jlab.org/entry/3777566
85
2/4 12-2pm6196-6200
bpm calibration through Compton ffb on, at 140, pita scan, detector cal, AposU/V scan, Hel mag scan.
applies to run5897 onwards, could be used on runs after new target installed 5853ish
https://logbooks.jlab.org/entry/3778029
86
2/5
laser macropulse chassis getting replaced, aperture 2 maybe will get removed from injector
87
88
2/4 6:55 pm slug131,132
flipper edited, slugs 31-out will have AposV-400 from nominal,and 32-in to have AposV-400 from nominal
89
2/5 9:39am slug132,133..
flipper edited, slugs 33-out will have AposV-800 from nominal,and 32-in to have AposV-400 from nominal
90
2/5 day sometime slug131, post run6220
ioc rtp ia down, flipper run so -800 AposV?
91
2/66242unser likecameron did
https://logbooks.jlab.org/entry/3786176
171.150.00945315867.978421.07858733.126513211
92
2/7new orbit tgt lock implimented
93
2/7 9:20am AT moved6271ATout moved in both arms
https://logbooks.jlab.org/entry/3779345
94
2/8-2/9Moller meas
https://logbooks.jlab.org/entry/3780396
95
2/10 ATout RHRS moved63372/10 ATout RHRS moved
https://logbooks.jlab.org/entry/3781210
96
2/10 morning (~11am)
slug =4000, ihwp in for transverse running (had been slug134,ihwpin)
97
2/10 11:31 AM
edit flipper to nominal values for IHWPout for transverse running
98
2/106348AposU/V scan
99
2/10/2020 18:46:00 slug4003 6351
edit flipper for IHWP-IN APosV, cause 4aX,0L05Y largish, takes effect for slug4003 (Ca40, IHWPin)
https://logbooks.jlab.org/entry/3781471
100
2/10/2020 11:25 pm for slug 4006 or 4007
edit flipper for IHWP-IN APosV, split difference between nominal and slug4004, so AposV=-735. will take effect for IHWP-in, next In slug (4006 or 4007)
https://logbooks.jlab.org/entry/3781556